# Policy Mirror Descent with Lookahead

### v1

@@ 5.1 Monte Carlo $h$-Greedy Policy Evaluation
Estimating the lookahead action value function $Q_{h}^{\pi}$ for a given policy $\pi$ involves in particular solving a $h$-horizon planning problem using samples from the MDP. Our estimation procedure combines a planning method with Monte Carlo estimation under a generative model of the MDP.
We now describe our approach to estimate $Q_{h}^{\pi}(s,a)$ at a given fixed state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$. We define a sequence $(\mathcal{S}_{k})$ of sets of visited states as follows: Set $\mathcal{S}_{h}=\{s\}$ and define for every $k\in[0,h-1]$, for every $\tilde{s}\in\mathcal{S}_{k+1}$ and for ever $\tilde{a}\in\mathcal{A}$ the multiset $\mathcal{S}_{k}(\tilde{s},\tilde{a}):=\{s_{j}^{\prime}\sim P(\cdot|\tilde{s},\tilde{a}):j\in[M_{k+1}]\}$ (including duplicated states) where $M_{k+1}$ is the number of states sampled from each state $\tilde{s}\in\mathcal{S}_{k+1}\,,$ and $\mathcal{S}_{k}=\bigcup_{\tilde{s}\in\mathcal{S},\tilde{a}\in\mathcal{A}}\mathcal{S}_{k}(\tilde{s},\tilde{a})\cup\mathcal{S}_{k+1}$ (without duplicated states).
The estimation procedure is recursive. First we define the value function $\hat{V}_{1}$ as follows:
$\hat{V}_{1}(\tilde{s}):=\hat{V}^{\pi}(\tilde{s})\,,\quad\forall\tilde{s}\in\mathcal{S}_{0}\,,$ (6)
where $\hat{V}^{\pi}(\tilde{s})$ is an estimate of $V^{\pi}(\tilde{s})$ computed using $M_{0}$ rollouts of the policy $\pi$ from state $\tilde{s}\,.$ More precisely, we sample $M_{0}$ trajectories $(\tau_{j}(\tilde{s}))_{j\in[M_{0}]}$ of length $H\geq 1$ rolling out policy $\pi$, i.e. $\tau_{j}(\tilde{s}):=\{(s_{k}^{j},a_{k}^{j},r(s_{k}^{j},a_{k}^{j}))\}_{0\leq k\leq H-1}$ with $s_{0}^{j}=\tilde{s}$ and $a_{k}^{j}\sim\pi(\cdot|s_{k}^{j})$. Then the Monte Carlo estimator is computed as $\hat{V}^{\pi}(\tilde{s})=\frac{1}{M_{0}}\sum_{j=1}^{M_{0}}\sum_{k=0}^{H-1}\gamma^{k}r(s_{k}^{j},a_{k}^{j})\,.$ For each stage $k\in\left[1,h\right]$ for $h\geq 1$, we define the following estimates for every $\tilde{s}\in\mathcal{S}_{k},\tilde{a}\in\mathcal{A}$,
$\hat{Q}_{k}(\tilde{s},\tilde{a})$ $:=r(\tilde{s},\tilde{a})+\frac{\gamma}{M_{k}}\sum_{s^{\prime}\in\mathcal{S}_{k-1}}\hat{V}_{k}(s^{\prime})\,,$
$\hat{V}_{k+1}(\tilde{s})$ $:=\max_{\tilde{a}\in\mathcal{A}}\hat{Q}_{k}(\tilde{s},\tilde{a})\,,$ (7)
@@
Notice that actions are exhaustively selected in our procedure like in a planning method. Using bandit ideas to guide Monte Carlo planning (see e.g. Kocsis and Szepesvári (2006a)) would certainly be interesting to investigate for more efficiency. We leave the investigation of such selective action sampling and exploration for future work. In practice, lookahead policies are often computed using tree search techniques (Efroni et al., 2018a).
@@ 5.2 Analysis of Inexact $h$-PMD
As can be directly seen from this formulation, the $\gamma^{h}$ convergence rate from the exact setting generalizes to the inexact setting. Thus, as before, faster rates of convergence are achieved for higher depths of lookahead. Another improvement with relation to Johnson et al. (2023) is that the additive error term has a weaker dependence on the effective horizon $1-\gamma$: the term $\frac{2b}{(1-\gamma)(1-\gamma^{h})}$ is smaller for all values of $h\geq 2$ than $\frac{4b}{(1-\gamma)^{2}}$. Therefore the algorithm converges to better solutions (in terms of suboptimality gap) as the asymptotic error on the suboptimality gap achievable by our algorithm (when $k\to\infty$) is smaller for deeper lookahead.
@@
Assume that inexact $h$-PMD is run for a number of iterations $K>\frac{1}{h(1-\gamma)}\log(\frac{4}{\epsilon(1-\gamma)(1-\gamma^{h})})$, using the choice of stepsize defined by the sequence $(c_{k}):=(\gamma^{2h(k+1)})$. Additionally, suppose we are given a target suboptimality value $\epsilon>0$, and a probability threshold $\delta>0$. Finally, assume the lookahead value function $\hat{Q}_{h}^{\pi_{k}}$ is approximated at each iteration with the Monte Carlo estimation procedure described in section 5.1, with the following parameter values:
 $\forall\,1\leq j\leq h\,,M_{j}=M>\frac{144}{(1-\gamma)^{6}\varepsilon^{2}}\log(4h|\mathcal{A}||\mathcal{S}_{0}||\mathcal{C}|/\delta)\,,$
$M_{0}=\frac{144\gamma^{2h}}{(1-\gamma)^{4}(1-\gamma^{h})^{2}\varepsilon^{2}}\log(4|\mathcal{S}_{0}||\mathcal{C}|/\delta)\,,$
Then, with probability at least $1-\delta$, the suboptimality at all iterations $k$ satisfy the following bound:
@@
Using this procedure, $h$-PMD uses at most $KM_{0}H|\mathcal{S}|+hM|\mathcal{S}||\mathcal{A}|$ samples in total. The overall sample complexity of the inexact $h$-PMD is then given by $\tilde{\mathcal{O}}(\frac{|\mathcal{S}|}{h\epsilon^{2}(1-\gamma)^{6}(1-\gamma^{h})^{2}}+\frac{|\mathcal{S}||\mathcal{A}|}{\epsilon^{2}(1-\gamma)^{7}})$ where the notation $\tilde{O}$ hides at most polylogarithmic factors.
Compared to Theorem 5.1 in Johnson et al. (2023), the dependence on the effective horizon improves by a factor of the order of $1/(1-\gamma)$ for $h>|\mathcal{A}|^{-1}(1-\gamma)^{-1}$ thanks to our Monte Carlo lookahead estimator. Their sample complexity is of order $\tilde{\mathcal{O}}(\frac{1}{\epsilon^{2}(1-\gamma)^{8}})$, whereas for $h>\frac{1}{|\mathcal{A}|(1-\gamma)}$ ours becomes $\tilde{\mathcal{O}}(\frac{1}{\epsilon^{2}(1-\gamma)^{7}})$. Similarly to Johnson et al. (2023), our sample complexity does not depend on quantities such as the distribution mismatch coefficient used in prior work (Xiao, 2022; Lan, 2023), which may scale with the state space size. We provide a proof of this theorem including more detailed lower bounds for the parameters in the appendix.
@@ 6 Extension to Function Approximation
In MDPs with prohibitively large state action spaces, we represent each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ by a feature vector $\psi(s,a)\in{{\mathbb{R}}}^{d}$ where typically $d\ll|\mathcal{S}|\times|\mathcal{A}|$ and where $\psi:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{d}$ is a feature map, also represented as a matrix $\Psi\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|\times d}$. Assuming that this feature map holds enough “information” about each state action pair (more rigorous assumptions will be provided later), it is possible to approximate any value function using a matrix vector product $\Psi\theta$, where $\theta\in\mathbb{R}^{d}$. Following this idea, we propose a novel generalization of inexact $h$-PMD with function approximation by using this new value function for the inexact $h$-PMD update.
@@ 6.1 Inexact $h$-PMD with Function Approximation
Using the above notations, assume access to $\theta_{k}$, which we use to approximate $Q^{\pi_{k}}_{h}$ as $\Psi\theta_{k}$. Then the inexact $h$-PMD update rule can be adapted as follows:
@@
We note that the policy update above can be implicit to avoid computing a policy for every state. It is possible to implement this algorithm using our Monte Carlo planning method in a similar fashion to Lattimore and Szepesvari (2019): use the procedure in 5.1 to estimate $\hat{Q}_{h}(s,a)$ for all state action pairs $(s,a)$ in some set $\mathcal{C}\subset\mathcal{S}\times\mathcal{A}$, and use these estimates as targets for a least squares regression to approximate $\hat{Q}_{h}$ by $\Psi\hat{\theta}$. We provide a specific procedure to accomplish this in the appendix. We conclude this section with a convergence analysis of $h$-PMD adapted to use linear function approximation.
@@ 6.2 Convergence Analysis of $h$-PMD with Linear Function Approximation
Our analysis follows the approach of Lattimore and Szepesvari (2019). We make the following standard assumptions on the feature map $\psi$, given by $\Psi$ in matrix form. We consider $h$ as a fixed input to the algorithm.
@@
The feature matrix $\Psi$ is full rank.
@@
There exists $\epsilon>0$ s.t. for any $\pi\in\Pi$, $\inf_{\theta\in\mathbb{R}^{d}}\lVert Q^{\pi}_{h}-\Psi\theta\rVert_{\infty}\leq\epsilon\,.$
@@
Let $(\theta_{k})$ be the sequence of iterates produced by $h$-PMD with linear function approximation, run for $K$ iterations, starting with policy $\pi_{0}\in\operatorname{ri}(\Pi)$, using stepsizes defined by $(c_{k})$. Assume that targets $\hat{Q}_{h}^{\pi_{k}}$ were computed using the procedure described in sec. 5.1 such that with probability at least $1-\delta$ for we have that $\forall k\;\leq K\,,\forall z\in\mathcal{C}\subset\mathcal{S}\times\mathcal{A}$ the targets satisfy $|\hat{Q}_{h}^{\pi_{k}}(z)-Q_{h}^{\pi_{k}}(z)|\leq\epsilon$ [^2] This can be achieved by choosing the parameters as in Theorem 5.4, with some slight modications presented in appendix C. Finally, for each iteration assume there exists some $\theta\in\mathbb{R}^{d}$ for which $\lVert Q^{\pi_{k}}_{h}-\Psi\theta\rVert_{\infty}\leq\lVert\epsilon_{\pi_{k}}\rVert_{\infty}$. There exists a choice of $\mathcal{C}$ for which the policies computed (or implicitly defined) at each iteration satisfy:
$\lVert V^{\star}-{V^{\pi_{k}}}\rVert_{\infty}$ $\leq\gamma^{hk}\left(\lVert V^{\star}-V^{\pi_{0}}\rVert_{\infty}+\frac{1}{1-\gamma}\sum_{t=1}^{k}\frac{c_{t-1}}{1-\gamma}\right)+\frac{2\max_{i\leq k}\lVert\epsilon_{\pi_{i}}\rVert_{\infty}(1+\sqrt{d})+2\sqrt{d}\,\epsilon}{(1-\gamma)(1-\gamma^{h})}\,.$
Notice that our performance bound does not depend on the state space size like in sec. 5 and depends instead on the dimension $d$ of the feature space. We should notice though that the choice of the set $\mathcal{C}$ is crucial for this result and given by the Kiefer-Wolfowitz theorem (Kiefer and Wolfowitz, 1960) which guarantees that $|\mathcal{C}|=\mathcal{O}(d^{2})$ as discussed in Lattimore and Szepesvari (2019).
As in the previous sections, our method for function approximation yields a sample complexity that is instance independent. In comparison to Alfano et al. (2023); Yuan et al. (2023), their results depend on distribution mismatch coefficients which can scale with the size of the state space.
Figure 1: Suboptimality value function gap for $h$-PMD in the exact (left) and inexact (right) settings. (Right) We performed 32 runs for each value of $h$, the mean is shown as a solid line and the standard deviation as a shaded area.
@@ 7 Simulations
We conduct simulations to investigate the effect of the lookahead depth on the convergence rate and illustrate our theoretical findings. We run the $h$-PMD algorithm for different values of $h$ in both exact and inexact settings on the DeepSea environment from DeepMind’s bsuite (Osband et al., 2020) using a grid size of 64 by 64, and a discount factor $\gamma=0.99$.
Exact $h$-PMD. We run the exact $h$-PMD algorithm for 100 iterations for increasing values of $h$ using the KL divergence. Similar results were observed for the Euclidean divergence.
We tested two different step size schedules: (a) in dotted lines in Fig. 1, $\eta_{k}$ equal to its lower bound in sec. 4, with the choice $c_{k}:=\gamma^{2h(k+1)}$ (note the dependence on $h$); and (b) in solid lines, $\eta_{k}$ identical stepsize schedule across all values of $h$ with $c_{k}:=\gamma^{2(k+1)}$ to isolate the effect of the lookahead. We clearly observe under both stepsize schedules that $h$-PMD converges in fewer iterations when lookahead value functions are used instead of regular $Q$ functions.
Inexact $h$-PMD. In this setting, we estimated the value function $Q_{h}$ using the vanilla Monte Carlo planning procedure detailed in section 5 in a stochastic variant of the DeepSea environment, and all the parameters except for $h$ were kept identical across runs. As predicted by our results, $h$-PMD converges in fewer iterations when $h$ is increased. We also observed in our simulations that $h$-PMD uses less samples overall (see Appendix D for the total number of samples used at each iteration), and usually converges after less overall computation time.

### v2

@@ 5.1 Monte Carlo $h$-Greedy Policy Evaluation
Estimating the lookahead action value function $Q_{h}^{\pi}$ for a given policy $\pi$ involves solving a $h$-horizon planning problem using samples from the MDP. Our estimation procedure combines a standard planning method with Monte Carlo estimation under a generative model of the MDP. We give an algorithmic description of the procedure below, and defer the reader to Appendix B.1 for a precise definition of the procedure. Applying the recursive algorithm below with $k:=h$ returns an estimate for the action value function $Q_{h}^{\pi}(s,a)$ at a given state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$.
Algorithm 1 Lookahead Q-function Estimation via Monte Carlo Planning
Procedure $Q(k,s,a,\pi)$
if $k=1$ then
     return $r(s,a)+\frac{\gamma}{M}\sum_{i=1}^{M}\hat{V}^{\pi}(s_{i}^{\prime})$ (where $s_{i}^{\prime}\sim\mathcal{P}(\cdot|s,a)$ for $i\in[M]$)
else
     return $r(s,a)+\frac{\gamma}{M}\sum_{i=1}^{M}\max_{a^{\prime}\in\mathcal{A}}Q(k-1,s_{i}^{\prime},a^{\prime},\pi)$ (where $s_{i}^{\prime}\sim\mathcal{P}(\cdot|s,a)$ for $i\in[M]$)
end if
Note that the base case of the recursion estimates the value function using Monte Carlo rollouts, see Appendix B.1.
@@
Notice that actions are exhaustively selected in our procedure like in a planning method. Using bandit ideas to guide Monte Carlo planning (see e.g. Kocsis and Szepesvári (2006a)) would be interesting to investigate for more efficiency. We leave the investigation of such selective action sampling and exploration for future work. In practice, lookahead policies are often computed using tree search techniques (Efroni et al., 2018a).
@@ 5.2 Analysis of Inexact $h$-PMD
As can be directly seen from this formulation, the $\gamma^{h}$ convergence rate from the exact setting generalizes to the inexact setting. Therefore, higher lookahead depths lead to faster convergence rates. Another improvement with relation to Johnson et al. (2023) is that the additive error term has a milder dependence on the effective horizon $1-\gamma$: the term $\frac{2b}{(1-\gamma)(1-\gamma^{h})}$ is smaller for all values of $h\geq 2$ than $\frac{4b}{(1-\gamma)^{2}}$. Larger lookahead depths yield a smaller asymptotic error in terms of the suboptimality gap.
@@
Assume that inexact $h$-PMD is run for a number of iterations $K>\frac{1}{h(1-\gamma)}\log(\frac{4}{\epsilon(1-\gamma)(1-\gamma^{h})})$, using the choice of stepsize defined by the sequence $(c_{k}):=(\gamma^{2h(k+1)})$. Additionally, suppose we are given a target suboptimality value $\epsilon>0$, and a probability threshold $\delta>0$. Finally, assume the lookahead value function $\hat{Q}_{h}^{\pi_{k}}$ is approximated at each iteration with the Monte Carlo estimation procedure described in section 5.1, with the following parameter values: $M_{0}=\tilde{\mathcal{O}}(\frac{\gamma^{2h}}{(1-\gamma)^{4}(1-\gamma^{h})^{2}\varepsilon^{2}})$ and for all $j\in[h]\,,M_{j}=M=\tilde{\mathcal{O}}(\frac{1}{(1-\gamma)^{6}\varepsilon^{2}})\,.$ Then, with probability at least $1-\delta$, the suboptimality at all iterations $k$ satisfy the following bound:
@@
Using this procedure, $h$-PMD uses at most $KM_{0}HS+hKMSA$ samples in total. The overall sample complexity of the inexact $h$-PMD is then given by $\tilde{\mathcal{O}}(\frac{S}{h\epsilon^{2}(1-\gamma)^{6}(1-\gamma^{h})^{2}}+\frac{SA}{\epsilon^{2}(1-\gamma)^{7}})$ where the notation $\tilde{\mathcal{O}}$ hides at most polylogarithmic factors.
Compared to Theorem 5.1 in Johnson et al. (2023), the dependence on the effective horizon improves by a factor of the order of $1/(1-\gamma)$ for $h>A^{-1}(1-\gamma)^{-1}$ thanks to our Monte Carlo lookahead estimator. Their sample complexity is of order $\tilde{\mathcal{O}}(\frac{1}{\epsilon^{2}(1-\gamma)^{8}})$, whereas for $h>\frac{1}{A(1-\gamma)}$ ours becomes $\tilde{\mathcal{O}}(\frac{1}{\epsilon^{2}(1-\gamma)^{7}})$.
Our sample complexity does not depend on quantities such as the distribution mismatch coefficient used in prior work (Xiao, 2022; Lan, 2023), which may scale with the state space size.
@@ 6 Extension to Function Approximation
In MDPs with prohibitively large state action spaces, we represent each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ by a feature vector $\psi(s,a)\in{{\mathbb{R}}}^{d}$ where typically $d\ll SA$ and where $\psi:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{d}$ is a feature map, also represented as a matrix $\Psi\in\mathbb{R}^{SA\times d}$. Assuming that this feature map holds enough “information” about each state action pair (more rigorous assumptions will be provided later), it is possible to approximate any value function using a matrix vector product $\Psi\theta$, where $\theta\in\mathbb{R}^{d}$. In this section, we propose an inexact $h$-PMD algorithm using action value function approximation.
@@ 6.1 Inexact $h$-PMD with Function Approximation
Using the notations above, approximating $Q^{\pi_{k}}_{h}$ by $\Psi\theta_{k}$, the inexact $h$-PMD update rule becomes:
@@
We note that the policy update above can be implicit to avoid computing a policy for every state. It is possible to implement this algorithm using our Monte Carlo planning method in a similar fashion to Lattimore and Szepesvari (2019): use the procedure in section 5.1 to estimate $Q_{h}^{\pi_{k}}(s,a)$ for all state action pairs $(s,a)$ in some set $\mathcal{C}\subset\mathcal{S}\times\mathcal{A}$, and use these estimates as targets for a least squares regression to approximate $Q_{h}^{\pi_{k}}$ by $\Psi\theta_{k}$ at time step $k$. See Appendix C for further details. We conclude this section with a convergence analysis of $h$-PMD with linear function approximation.
@@ 6.2 Convergence Analysis of $h$-PMD with Linear Function Approximation
Our analysis follows the approach of Lattimore and Szepesvari (2019). We make the following standard assumptions.
@@
The feature matrix $\Psi\in{{\mathbb{R}}}^{SA\times d}$ where $d\leq SA$ is full rank.
@@
There exists $\epsilon_{\text{FA}}>0$ s.t. for any $\pi\in\Pi$, $\inf_{\theta\in\mathbb{R}^{d}}\lVert Q^{\pi}_{h}-\Psi\theta\rVert_{\infty}\leq\epsilon_{\text{FA}}\,.$
We defer a discussion about Assumption 6.2 to Appendix C.2.
@@
Let $(\theta_{k})$ be the sequence of iterates produced by $h$-PMD with linear function approximation, run for $K$ iterations, starting with policy $\pi_{0}\in\operatorname{ri}(\Pi)$, using stepsizes defined by $(c_{k})$. Assume that targets $\hat{Q}_{h}^{\pi_{k}}$ were computed using the procedure described in sec. 5.1 such that with probability at least $1-\delta,\,$ $\forall k\;\leq K\,,\forall z\in\mathcal{C}\subset\mathcal{S}\times\mathcal{A}$ the targets satisfy $|\hat{Q}_{h}^{\pi_{k}}(z)-Q_{h}^{\pi_{k}}(z)|\leq\epsilon$ [^2] This can be achieved by choosing the parameters as in Theorem 5.4 (see Appendix C for details).
Then, there exists a choice of $\mathcal{C}$ for which the policy iterates satisfy at each iteration:
$\lVert V^{\star}-{V^{\pi_{k}}}\rVert_{\infty}$ $\leq\gamma^{hk}\left(\lVert V^{\star}-V^{\pi_{0}}\rVert_{\infty}+\frac{1}{1-\gamma}\sum_{t=1}^{k}\frac{c_{t-1}}{1-\gamma}\right)+\frac{2\sqrt{d}\,\epsilon+2\,(1+\sqrt{d})\epsilon_{\text{FA}}}{(1-\gamma)(1-\gamma^{h})}\,.$
The proof of Theorem 6.3 can be found in Appendix C.3. Notice that our performance bound does not depend on the state space size like in sec. 5 and depends on the dimension $d$ of the feature space instead. We should notice though that the choice of the set $\mathcal{C}$ is crucial for this result and given by the Kiefer-Wolfowitz theorem (Kiefer and Wolfowitz, 1960) which guarantees that $|\mathcal{C}|=\mathcal{O}(d^{2})$ as discussed in Lattimore and Szepesvari (2019). Moreover, our method for function approximation yields a sample complexity that is instance independent. Existing results for PMD with function approximation Alfano et al. (2023); Yuan et al. (2023) depend on distribution mismatch coefficients which can scale with the size of the state space.
In terms of computational complexity, at each iteration, $h$-PMD uses $O((MA)^{h})$ elementary operations where $M$ is the size of the minibatch of trajectories used, $A$ is the size of the action space and $h$ is the lookahead depth. This computational complexity which scales exponentially in $h$ is inherent to tree search methods. Despite this seemingly prohibitive computational scaling, tree search has been instrumental in practice for achieving state of the art results in some environments (Schrittwieser et al., 2019; Hessel et al., 2022) and modern implementations (notably using GPU-based search on vectorized environments) make tree search much more efficient in practice.
Figure 1: Suboptimality value function gap for $h$-PMD in the exact (left) and inexact (middle/right) settings, plotted against iterations in the exact case (left) and against both iterations (middle) and runtime (right) in the inexact case. 16 runs performed for each $h$, mean in solid line and standard deviation as shaded area. In dotted lines (left), the step size $\eta_{k}$ is equal to its lower bound in sec. 4, with the choice $c_{k}:=\gamma^{2h(k+1)}$ (note the dependence on $h$) and in solid lines, the step size $\eta_{k}$ is set using an identical stepsize schedule across all values of $h$ with $c_{k}:=\gamma^{2(k+1)}$ to isolate the effect of the lookahead. Notice that higher values of $h$ still perform better even in terms of runtime.
@@ 7 Simulations
We conduct simulations to investigate the effect of the lookahead depth on the convergence rate and illustrate our theoretical findings. We run the $h$-PMD algorithm for different values of $h$ in both exact and inexact settings on the DeepSea environment from DeepMind’s bsuite (Osband et al., 2020) using a grid size of 64 by 64, and a discount factor $\gamma=0.99$. Additional experiments are provided in Appendix D. Our codebase where all our experiments can be replicated is available here: https://gitlab.com/kimon.protopapa/pmd-lookahead.
Exact $h$-PMD. We run the exact $h$-PMD algorithm for 100 iterations for increasing values of $h$ using the KL divergence. Similar results were observed for the Euclidean divergence. We tested two different stepsize schedules: (a) in dotted lines in Fig. 1 (left), $\eta_{k}$ equal to its lower bound in sec. 4, with the choice $c_{k}:=\gamma^{2h(k+1)}$ (note the dependence on $h$); and (b) in solid lines, $\eta_{k}$ identical stepsize schedule across all values of $h$ with $c_{k}:=\gamma^{2(k+1)}$ to isolate the effect of the lookahead. We clearly observe under both stepsize schedules that $h$-PMD converges in fewer iterations when lookahead value functions are used instead of regular $Q$ functions.
Inexact $h$-PMD. In this setting, we estimated the value function $Q_{h}$ using the vanilla Monte Carlo planning procedure detailed in section 5 in a stochastic variant of the DeepSea environment, and all the parameters except for $h$ were kept identical across runs. As predicted by our results, $h$-PMD converges in fewer iterations when $h$ is increased (see Fig. 1 (middle)). We also observed in our simulations that $h$-PMD uses less samples overall (see Appendix D for the total number of samples used at each iteration), and usually converges after less overall computation time (see Fig. 1 (right)). We also refer the reader to Appendix D where we have performed additional experiments with a larger lookahead depth $h=100.$ In this case, the algorithm converges in a single iteration. This is theoretically expected as computing the lookahead values with very large $h$ boils down to computing the optimal values like in value iteration with a large number of iterations. We also performed additional experiments in continuous control tasks to illustrate the general applicability of our algorithm (see Fig. 8 in Appendix D).
Remark 7.1.
(Choice of the lookahead depth $h$.) The lookahead depth is a hyperparameter of the algorithm and can be tuned similarly to other hyperparameters such as the step size of the algorithm. Of course, the value would depend on the environment and the structure of the reward at hand. Sparse and delayed reward settings will likely benefit from lookahead with larger depth values. We have performed several simulations with different values of $h$ for each environment setting and the performance can potentially improve drastically with a better lookahead depth value (see also appendix D for further simulations). In addition, we observe that larger lookahead depth is not always better: see Fig. 7 in the appendix for an example where large lookahead depth becomes slower and does not perform better. Note that in this more challenging practical setting the best performance is not obtained for higher values of $h$: intermediate values of $h$ perform better. This illustrates the tradeoff in choosing the depth $h$ between an improved convergence rate and the computational cost induced by a larger $h$. We believe further investigation regarding the selection of the depth parameter might be useful to further improve the practical performance of the algorithm.